OpenStudy (anonymous):

1. The half-life of a radioactive substance is 5 minutes. how much of an original mass of 100 grams remains after half an hour? 2. What is the area of triangle ABC if a=11, b=13 and c=31 degrees 10 3.Two sides of a triangular plot of land have the lengths of 400 ft and 600 ft. The measurement of the angle between those sides is 46 degrees 20.Find the perimeter of the plot.

OpenStudy (anonymous):

plzzzz help me

OpenStudy (anonymous):

i don't know the formula?

OpenStudy (stamp):

\[1.\ N(t)=N _{0}e^{kt}\]\[N_{0}=100 \ g\]

OpenStudy (stamp):

t = 1/2 hour k = 5/60 hour = 1 / 12 hour

OpenStudy (stamp):

It is a decay problem so I think you would use k = -/12 hour

OpenStudy (anonymous):

-12?

OpenStudy (anonymous):

k=-12?

OpenStudy (stamp):

I am a bit confused on the wording in 2 and 3. Can you double check and make sure they are accurate? Also, - 1 / 12 hour. I am still reviewing half-lifes to make sure I give you the correct solution.

OpenStudy (anonymous):

why is it 1/2 from 5/60 it is half an hour so 30 minutes. . . why do you use 60? please explain. . .

OpenStudy (stamp):

\[ Halflife = \frac {-ln(2)}{k} \]\[k=rate \ of \ decay\]\[ Halflife =30 \ min\]\[ 30 \ min = \frac{-ln2}{k} \]\[ k = \frac{-ln2}{30\ min} \]\[ t =5\ min \]

OpenStudy (anonymous):

hmhm ok. . can we proceed to the second question

OpenStudy (stamp):

\[N_{0}=x \]\[ N = 100\ g\]\[k=\frac{-ln2}{30\ min}\]\[t=5\ min\] \[x = 100e^{5\frac{-ln2}{30}} \]\[x=89\ g\]I wouldn't mind getting a second take on this answer

OpenStudy (stamp):

I would like to do 2 and 3, but they still seem off. 

OpenStudy (anonymous):

c=31degrees10

OpenStudy (stamp):

Hmm, that still confuses me. Is this the correct visualisation for 3?

OpenStudy (anonymous):

then. . .

OpenStudy (anonymous):

that is 46 degrees 10

OpenStudy (stamp):

46°10 does not make sense...

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

how to solve it what will we use ? trigonometric functions

OpenStudy (stamp):

We could use trigonometry if we know it is a right triangle. The problem descriptions still seem incomplete.

OpenStudy (stamp):

Law of Sine or Law of Cosines? are these topics you have recently covered in lecture?

OpenStudy (anonymous):

a little bit

OpenStudy (anonymous):

hhhhhhhhhhhhheeeeeeeeeeeeeellllllllllppppppppp

OpenStudy (stamp):

I really do not think we will be able to help you. I repeat, problems 2 and 3 simply do not make sense. The metrics provided are incoherent.

OpenStudy (stamp):

What class title is this? High school, college?

OpenStudy (anonymous):

higgh school

OpenStudy (stamp):

Algebra I/II/pre, ... Precalculus...

OpenStudy (anonymous):

pre

OpenStudy (anonymous):

stamp?

OpenStudy (stamp):

Are these problems coming from a textbook, a worksheet, or some other source

OpenStudy (anonymous):

Its for National Achievement Test

OpenStudy (stamp):

Retype Problem 2. Just problem 2. Nothing else except Problem 2, word for word.

OpenStudy (anonymous):

What is the area of \[\Delta ABC\]

OpenStudy (anonymous):

if a=11, b=13 and C=31DEGREES 10

OpenStudy (anonymous):

\[31^{\sigma}10\]